OpenStudy (anonymous):

14+6sqrt5=? a)2+sqrt5 b) 3+sqrt5 c) 5+sqrt3 d)3+2sqrt5

OpenStudy (anonymous):

Something is wrong. None of these answers are right.

OpenStudy (kc_kennylau):

Let 14+6sqrt5 be (a+sqrtb)^2

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} 14+6\sqrt5&=&(a+\sqrt b)^2\\ 14+6\sqrt5&=&a^2+b+2a\sqrt b\\ a^2+b=14&&2a\sqrt b=6\sqrt5 \end{array}\]

OpenStudy (kc_kennylau):

\[4a^2b=180\]\[a^2b=45\]

OpenStudy (kc_kennylau):

a^2=9 & b=5

OpenStudy (kc_kennylau):

Therefore it's the option b

OpenStudy (kc_kennylau):

CHECK:\[(3+\sqrt5)^2=9+2\cdot3\sqrt5+5=14+6\sqrt5\]